Todd currently has $5,000. What was the value of his money four years ago if he's earned
4 percent interest each year?

Answer by stanbon(75887)

Todd currently has $5,000. What was the value of his money four years ago if he's earned
4 percent interest each year?
----
A(t) = P(1+rt)
----
5000 = P(1 + 0.04*4)
---
5000 = P(1.16)
P = 4310.34
=================
Cheers,
Stan H.